# Improving causal inference in interrupted time series analysis: the triple difference design

### 18 Mar 2026 v1

@@ Abstract
Background: Interrupted time series analysis (ITSA) is a valuable quasi-experimental design for evaluating health policy and intervention effects. While multiple-group ITSA (MG-ITSA) strengthens causal inference by incorporating a control group, residual confounding from unmeasured time-varying factors may persist. The triple-difference interrupted time series (DDD-ITSA) design extends this framework by adding a second control group to further isolate treatment effects, yet the design remains underutilized and lacks formal methodological guidance.
@@
Conclusions: The DDD-ITSA design offers a robust approach for strengthening causal inferences when two-group comparisons may be confounded. By leveraging an additional control group, researchers can difference out remaining biases and assess effect heterogeneity across subgroups. The updated itsa package for Stata facilitates implementation, making this design accessible for routine application in health policy research. Careful attention to control group selection, baseline balance, and autocorrelation remains essential for valid interpretation.
@@
difference-in-difference-in-differences, triple differences, interrupted time series analysis, controlled interrupted time series analysis, multiple-group interrupted time series analysis
@@ 1 Background
Multiple-group or controlled ITSA (MG-ITSA) addresses these threats by introducing control units that were not exposed to the intervention but share similar baseline levels, pre-intervention trends, and observed characteristics with the treated unit [14, 1, 18]. By differencing outcomes between treated and control series over time, MG-ITSA relaxes the strong assumption of stable counterfactual trends within a single unit and instead relies on a parallel-trends-type assumption across groups [14, 18]. Under this framework, causal effects are identified by removing secular trends and common shocks that affect all units similarly, aligning MG-ITSA conceptually with difference-in-differences estimators [23]. However, when treatment effects are expected to vary across an additional dimension, such as subpopulations, exposure intensity, or policy eligibility, standard MG-ITSA may remain insufficient to isolate the causal effect of interest.
@@
A triple-difference interrupted time series (DDD-ITSA) design extends the logic of the MG-ITSA by adding a second control group for comparison. This approach mirrors the difference-in-differences-in-differences (DDD) framework in the econometrics literature, where an additional difference is used to remove remaining confounding that may persist after a standard two-group comparison [21]. In the ITSA context, the triple-difference estimator captures whether the post-intervention change in level and/or trend for the treated group is larger (or smaller) for the subgroup of interest than would be expected based on concurrent changes in control groups and comparison subgroups. By differencing across time, groups, and subgroups simultaneously, the DDD-ITSA design strengthens causal identification under assumptions analogous to those in the difference-in-differences (DiD) literature, namely that, absent the intervention, these differences would have evolved similarly over time [23].
@@
Applications that explicitly implement a DDD-ITSA framework remain relatively uncommon in the literature. One applied example is the evaluation of Maryland’s global budget revenue (GBR) model, where researchers used an interrupted time-series design with difference-in-differences comparisons and a triple differences analysis to assess changes in emergency department utilization, admissions, and revisits following GBR implementation, including how these trends varied across racial/ethnic and payer groups [7]. In this study, the triple-difference component helped identify differential effects of the policy on emergency department return outcomes across subpopulations beyond general statewide trends and group-level effects. Another example is research on the effects of the Affordable Care Act (ACA) on health insurance coverage among Latino subgroups, which combined ITSA with triple-difference models to evaluate heterogeneous changes in uninsured rates associated with the ACA and Medicaid expansion across multiple Latino ethnic groups [8].
@@ 2.1 The DDD-ITSA design
The DDD-ITSA design is an extension of the MG-ITSA design to an additional control group that is either geographically distant from the treatment and primary control groups, or comprises a different demographic that still allows for comparison (e.g., type of insurance coverage, different medical condition, ethnicity, etc.). This second control group may be thought of as adding a “doubly robustness” property to the analysis, in that it provides a second opportunity to get the causal inference correct if the treatment versus primary control is confounded. Figure 1 illustrates four of the many potential outcome scenarios that we could anticipate in the DDD-ITSA design when emphasizing differences in trends: (A) a treatment effect in the primary analysis and secondary analysis; (B) no treatment effect in either the primary or secondary analysis; (C) a treatment effect in the secondary analysis but not in the primary analysis (evidence of confounding); and (D) a treatment effect in the primary analysis but not in the secondary analysis (suggesting external influences or spillover affecting the secondary control).
Figure 1: Four general outcome scenarios for the DDD-ITSA (A) treatment effect in both primary and secondary analyses, (B) no treatment effect, (C) treatment effect in secondary analysis only, (D) treatment effect in primary only.
@@ 2.2 The DDD-ITSA model
where the autocorrelation parameter $\rho$ is the correlation coefficient between adjacent error terms such that $|\rho|<1$, and the disturbances $u_{t}$ are independent $N(0,\sigma^{2})$ [10].
Table 1 presents the coefficients and linear combination terms for the pre-treatment and post-treatment trends of each of the three groups, and Table 2 presents the difference-in-differences of trend comparisons between groups.
@@
Tables 3 and 4 present the analogous coefficients and linear combination terms for the pre-treatment and post-treatment level changes of each of the three groups and their respective difference in differences.
@@ 2.3 Assessing balance on baseline level and trend
In non-experimental settings, it is critical to evaluate whether the treatment and control groups are balanced in both the level and slope of the outcome variable prior to the intervention. Balance across groups in baseline levels and trends can be formally assessed using the coefficients from Equation (1). Specifically, statistically insignificant estimates of $\beta_{4}$ and $\beta_{5}$ indicate no systematic differences between the treatment group and the primary control group in pre-intervention levels and trends, respectively. Likewise, statistical insignificance of the linear combinations $(\beta_{4}-\beta_{8})$ and $(\beta_{5}-\beta_{9})$ suggests baseline equivalence between the treatment group and the secondary control group in levels and trends. Finally, statistically insignificant $\beta_{8}$ and $\beta_{9}$ imply balance between the two control groups in pre-treatment levels and trajectories.
@@ 2.4 Example
The following example illustrates the analysis of the DDD-ITSA design using data from California’s 1988 enactment of Proposition 99, a voter-approved initiative that increased the cigarette excise tax by \$0.25 per pack and funded statewide anti-smoking efforts. The outcome examined is per-capita cigarette sales (packs), measured annually at the state level from 1970 to 2000, with 1989 marking the first post-intervention year. The dataset includes cigarette sales and four covariates: average retail cigarette price, logged per-capita personal income, per-capita beer consumption, and the share of the population aged 15–24.
For exposition purposes, Idaho and Montana are assigned to the primary control group and Colorado as the secondary control state. These states were identified as balanced matches to California (the treated state) on baseline level and trend of cigarette sales using the itsamatch program in Stata [1]. We analyze the model using the DDD-ITSA feature in the itsa program for Stata [14]. Regression with Newey–West standard errors is specified, with a single lag autoregressive structure (AR1). The Stata post-estimation command lincom is used to compute Wald estimates for linear combinations of coefficients. The exact code used in this example is provided in the Appendix.
@@ 3 Results
Table 5 presents the regression results for the DDD-ITSA model. First, we review how well the groups are balanced on baseline level and trend. $\beta_{4}$ and $\beta_{5}$ indicate that California and the primary control states are balanced on baseline level and trend. Linear combinations indicate that California and Colorado are also balanced. Finally, $\beta_{8}$ and $\beta_{9}$ indicate that the two control groups are balanced.
Next, we evaluate treatment effects. $\beta_{7}$ indicates that California experienced a statistically significant annual net reduction of -2.07 per-capita cigarette sales compared to the primary control states. Similarly, the linear combination $(\beta_{7}-\beta_{11})$ indicates a statistically significant annual net decrease relative to the secondary control. The $\beta_{11}$ coefficient indicates that the difference-in-differences between the two control groups was non-statistically significant.
@@
Coefficient Estimate Std Err Z P 95% LCL 95% UCL
$\beta_{0}$ $126.40$ $4.58$ $27.57$ $0.001$ $117.41$ $135.38$
$\beta_{1}$ 1.43-1.43 $0.43$ 3.35-3.35 $0.001$ 2.27-2.27 0.59-0.59
$\beta_{2}$ 11.43-11.43 $4.46$ 2.56-2.56 $0.010$ 20.17-20.17 2.68-2.68
$\beta_{3}$ $0.58$ $0.61$ $0.95$ $0.345$ 0.62-0.62 $1.78$
$\beta_{4}$ $5.83$ $6.20$ $0.94$ $0.347$ 6.33-6.33 $17.98$
$\beta_{5}$ 0.35-0.35 $0.57$ 0.61-0.61 $0.539$ 1.47-1.47 $0.77$
$\beta_{6}$ 8.63-8.63 $6.44$ 1.34-1.34 $0.180$ 21.25-21.25 $3.98$
$\beta_{7}$ 2.07-2.07 $0.75$ 2.77-2.77 $0.006$ 3.54-3.54 0.61-0.61
$\beta_{8}$ $12.29$ $6.84$ $1.80$ $0.072$ 1.11-1.11 $25.70$
$\beta_{9}$ 0.10-0.10 $0.69$ 0.15-0.15 $0.880$ 1.46-1.46 $1.25$
$\beta_{10}$ 6.48-6.48 $7.75$ 0.84-0.84 $0.403$ 21.67-21.67 $8.70$
$\beta_{11}$ 0.31-0.31 $0.88$ 0.35-0.35 $0.723$ 2.03-2.03 $1.41$
  • Note: Standard errors are Newey–West AR(1) adjusted.
@@ 4 Discussion
This paper formalizes the triple-difference interrupted time series (DDD-ITSA) design, an extension of the multiple-group ITSA framework that incorporates an additional control group to strengthen causal identification in health policy and intervention research. The DDD-ITSA approach combines the structural time-series logic of ITSA with the econometric rigor of difference-in-differences-in-differences estimators, offering researchers a powerful tool for estimating causal effects when standard two-group comparisons may remain confounded by unmeasured time-varying factors. The key contributions of this paper are the explicit articulation of the DDD-ITSA model, including clear parameter interpretation, and the worked example to illustrate the practical application of the DDD-ITSA design.
Researchers implementing DDD-ITSA face several issues with the design that require careful consideration. First, the choice of primary and secondary control groups is substantively important and should be guided by content expertise and empirical balance assessments. The primary control should be as similar as possible to the treatment unit on observed characteristics and pre-treatment trends [14, 18], while the secondary control provides an additional layer of adjustment, ideally one that shares the primary control’s susceptibility to common shocks but differs in ways that help isolate treatment effects.
@@
Third, researchers should review autocorrelation and partial autocorrelation functions [3] and formally test for autocorrelation using appropriate diagnostics (e.g., the actest command in Stata [2]) and specify error structures accordingly, whether through Newey–West standard errors [20], Prais–Winsten regression [22], Cochrane–Orcutt regression [5], instrumental variables [11], or other time-series estimators [9, 6].
Fourth, researchers should endeavor to collect as long of time series data as possible. Linden [19] found that in the MG-ITSA context, a longer study period provides the necessary statistical resolution to detect trend-based effects and mitigates the bias introduced by autocorrelation, which can otherwise severely undermine power. Researchers should therefore design DDD-ITSA studies with the expectation that detecting a meaningful trend change will demand a longer time horizon than simpler designs might require.
@@
The DDD-ITSA design is particularly well-suited to several common scenarios in health services and policy research. First, when researchers suspect that two-group comparisons may be confounded by concurrent policy changes, economic shocks, or secular trends that differentially affect subgroups, the addition of a second control group can difference out these common sources of bias. Second, DDD-ITSA is valuable when treatment effects are expected to vary across an additional dimension such as demographic subgroups, payer types, or clinical conditions. Third, the design is appropriate when researchers have access to multiple potential control groups but are uncertain which provides the most valid counterfactual.

### 25 Apr 2026 v2

@@ Abstract
Background: Interrupted time series analysis (ITSA) is a valuable quasi-experimental design for evaluating health policy and intervention effects. While multiple-group ITSA (MG-ITSA) can improve causal inference under additional identifying assumptions by incorporating a control group, residual confounding from unmeasured time-varying factors may persist. The triple-difference interrupted time series (DDD-ITSA) design extends this framework by adding a second control group to further isolate treatment effects, yet the design remains underutilized and lacks formal methodological guidance.
@@
Conclusions: The DDD-ITSA design offers a useful extension of the interrupted time series framework that can improve causal inference when its identifying assumptions are plausible. By leveraging an additional control group, researchers can difference out remaining biases and assess effect heterogeneity across groups or comparison dimensions. The updated itsa package for Stata facilitates implementation, making this design accessible for routine application in health policy research. Careful attention to control group selection, baseline balance, and autocorrelation remains essential for valid interpretation.
@@
triple differences, interrupted time series, difference-in-differences, causal inference, quasi-experimental design, health policy
@@ 1 Background
Multiple-group or controlled ITSA (MG-ITSA) addresses these threats by introducing control units that were not exposed to the intervention but share similar baseline levels, pre-intervention trends, and observed characteristics with the treated unit [7, 8, 9]. By differencing outcomes between treated and control series over time, MG-ITSA relaxes the strong assumption of stable counterfactual trends within a single unit and instead relies on a parallel-trends-type assumption across groups [7, 9]. Under this framework, causal effects are identified by removing secular trends and common shocks that affect all units similarly, aligning MG-ITSA with difference-in-differences estimators, which it generalizes to settings with multiple time points and allows for estimation of both level and trend changes. However, when treatment effects are expected to vary across an additional dimension (such as subpopulations, exposure intensity, policy eligibility, or across alternative comparison groups) standard MG-ITSA may be insufficient to isolate the causal effect of interest.
@@
A triple-difference interrupted time series (DDD-ITSA) design extends the logic of the MG-ITSA by adding a second control group for comparison. This approach mirrors the difference-in-differences-in-differences (DDD) framework in the econometrics literature, where an additional difference is used to remove remaining confounding that may persist after a standard two-group comparison [10]. In the ITSA context, the triple-difference estimator captures whether the post-intervention change in level and/or trend for the treated group differs along an additional dimension of comparison (such as between groups or comparison dimensions within a population or relative to an additional control group) beyond what would be expected based on concurrent changes in the primary comparison groups. By differencing across time, groups, and an additional comparison dimension, the DDD-ITSA design can improve causal inference under assumptions analogous to those in the difference-in-differences (DiD) literature, namely that, absent the intervention, these relative differences would have evolved similarly over time [11].
@@
Applications that explicitly implement a DDD-ITSA framework remain relatively uncommon in the literature. One applied example is the evaluation of Maryland’s global budget revenue (GBR) model, where researchers used an interrupted time-series design with difference-in-differences comparisons and a triple differences analysis to assess changes in emergency department utilization, admissions, and revisits following GBR implementation, including how these trends varied across racial/ethnic and payer groups [12]. In this study, the triple-difference component helped identify differential effects of the policy on emergency department return outcomes across subpopulations beyond general statewide trends and group-level effects. Another example is research on the effects of the Affordable Care Act (ACA) on health insurance coverage among Latino subgroups, which combined ITSA with triple-difference models to evaluate heterogeneous changes in uninsured rates associated with the ACA and Medicaid expansion across multiple Latino ethnic groups, highlighting substantial variation in coverage gains that would not be captured by simpler two-group comparisons [13]. More generally, the triple-difference framework identifies differences in treatment effects across groups (i.e., effect heterogeneity), while estimation of the treatment effect within a specific subgroup represents a special case that requires the additional assumption that the comparison group experiences no treatment effect.
@@ 2.1 The DDD-ITSA design
The DDD-ITSA design is an extension of the MG-ITSA design to an additional control group that is either geographically distant from the treatment and primary control groups, or comprises a different demographic that still allows for comparison (e.g., type of insurance coverage, different medical condition, ethnicity, etc.). This second control group provides an additional comparison that can help assess whether estimated effects are sensitive to alternative control group constructions. Figure 1 illustrates four of the many potential outcome scenarios that we could anticipate in the DDD-ITSA design when emphasizing differences in trends: (A) a treatment effect in the primary analysis and secondary analysis (the treatment unit’s post treatment trend increases, while the primary control group and secondary control groups’ post-treatment trends remain consistent with how they were trending in the pre-treatment period); (B) no treatment effect in either the primary or secondary analysis (all groups’ post-treatment trends continue trending consistently with how they were trending in the pre-treatment period); (C) a treatment effect in the secondary analysis but not in the primary analysis (there is evidence of confounding either by spill-over or by an external event); and (D) a treatment effect in the primary analysis but not in the secondary analysis (suggesting that something outside the intervention is influencing outcomes in the secondary control group such as an economic trend, seasonal effect, or spillover from a neighboring intervention). While Figure 1 focuses on stylized outcome scenarios, it also provides intuition for the underlying triple-difference logic by illustrating how comparisons across time, groups, and an additional comparison dimension can reveal whether observed changes are attributable to the intervention or to confounding influences.
Figure 1: Four general outcome scenarios for the DDD-ITSA. (A) A treatment effect in the primary analysis, (B) No treatment effect, (C) A treatment effect in the secondary analysis (evidence of confounding in the primary comparison), (D) A treatment effect in the primary analysis, but evidence of confounding in the secondary analysis.
2.2 Identification assumptions
The validity of the DDD-ITSA estimator rests on a set of identifying assumptions that extend the logic of difference-in-differences estimators [10]. Specifically, identification requires that, in the absence of the intervention, the difference between the two-component difference-in-differences comparisons would have remained constant over time. This condition can be understood as a generalized parallel trends assumption [11]. While the DDD estimator is sometimes described as requiring only a single parallel trends condition, in practice it implies that any bias affecting the primary comparison between the treatment and first control group must also affect the comparison between the second control group and the first control group in an equivalent way. Under this condition, the shared bias component is differenced out, yielding an unbiased estimate of the treatment effect [10]. If the sources of non-parallel trends differ across these comparisons, then the triple-difference estimator may remain biased. As such, careful selection of control groups and substantive justification of the identifying assumptions are critical for valid causal interpretation. Conceptually, the DDD-ITSA estimator can be understood as the difference between two difference-in-differences comparisons, with the additional comparison dimension providing a second contrast that allows shared biases to be differenced out. This perspective provides a simple conceptual interpretation of the design without requiring additional graphical representation.
@@ 2.3 The DDD-ITSA model
Although Equation (1) provides a convenient parameterization of group-level trends and intervention effects, it is important to distinguish between the statistical model and the assumptions required for causal identification. The regression specification describes the structure of the outcome over time, including baseline levels, trends, and their changes following the intervention. However, the ability to interpret the estimated coefficients causally does not arise from the functional form of the model itself, but rather from the validity of the identifying assumptions described above. In particular, causal interpretation relies on the comparability of trends across groups and the assumption that any bias affecting the component comparisons is equivalent and can be differenced out. The model therefore serves as a tool for estimation, while identification is grounded in the study design and underlying assumptions.
@@
where the autocorrelation parameter $\rho$ is the correlation coefficient between adjacent error terms such that $|\rho|<1$, and the disturbances $u_{t}$ are independent $N(0,\sigma^{2})$ (see [15] for a detailed discussion of autocorrelation in time-series regression models).
Table 1 presents the coefficients and linear combination terms for the pre-treatment and post-treatment trends of each of the three groups, and Table 2 presents the difference-in-differences of trend comparisons between groups. As shown, $(\beta_{7}-\beta_{11})$ represents the difference-in-differences in trends between the treated unit and the second control group. Due to the specified parameterization, this expression is algebraically equivalent to the triple-difference estimand, which can also be written as $(\beta_{7}-\beta_{11})=[(\beta_{3}+\beta_{7})-\beta_{3}]-[(\beta_{3}+\beta_{11})-\beta_{3}]$. This equivalence arises from the choice of reference group and the structure of the regression model, which causes the primary control group terms to cancel out. Importantly, this algebraic simplification reflects how the estimand is represented within the model and does not, by itself, justify a causal interpretation. Causal interpretation instead depends on the identifying assumptions described above, particularly the requirement that any bias affecting the component comparisons is equivalent and can be differenced out.
@@
Tables 3 and 4 present the analogous coefficients and linear combination terms for the pre-treatment and post-treatment level changes of each of the three groups and their respective difference in differences. The DDD estimate for the immediate level effect (that is, the effect occurring in the period immediately following introduction of the treatment) is given by $(\beta_{6}-\beta_{10})$, representing the additional post-intervention level change for the treatment group relative to the second comparison group, net of secular changes observed in the holdout. Together, these parameters quantify both the short-run (level) and longer-run (trend) causal effects of the intervention within a triple-difference interrupted time series framework.
@@ 2.4 Assessing balance on baseline level and trend
In non-experimental settings, it is critical to evaluate whether the treatment and control groups are balanced in both the level and slope of the outcome variable prior to the intervention. In a randomized controlled trial, the expectation is that baseline levels and trends are statistically equivalent across groups, supporting identification of causal effects. In observational studies, however, pre-treatment differences in intercepts or trends may violate the parallel trends assumption, potentially biasing estimates and complicating causal inference regarding the impact of the intervention on the outcome.
Balance across groups in baseline levels and trends can be formally assessed using the coefficients from Equation (1). Specifically, statistically insignificant estimates of $\beta_{4}$ and $\beta_{5}$ indicate no systematic differences between the treatment group and the primary control group in pre-intervention levels and trends, respectively. Likewise, statistical insignificance of the linear combinations $(\beta_{4}-\beta_{8})$ and $(\beta_{5}-\beta_{9})$ suggests baseline equivalence between the treatment group and the secondary control group in levels and trends. Finally, statistically insignificant $\beta_{8}$ and $\beta_{9}$ imply balance between the two control groups in pre-treatment levels and trajectories.
@@ 2.5 Example
The following example illustrates the analysis of the DDD-ITSA design using data from California’s 1988 enactment of Proposition 99, a voter-approved initiative that increased the cigarette excise tax by \$0.25 per pack and funded statewide anti-smoking efforts. The outcome examined is per-capita cigarette sales (packs), a standard proxy for smoking prevalence in the tobacco control literature, measured annually at the state level from 1970 to 2000, with 1989 marking the first post-intervention year. The dataset, compiled by Abadie et al. [8], includes cigarette sales and four covariates: average retail cigarette price, logged per-capita personal income, per-capita beer consumption, and the share of the population aged 15–24. Eleven states that implemented major tobacco control programs during 1989–2000 were excluded, leaving 38 states as potential controls.
For exposition purposes, Idaho and Montana are assigned to the primary control group and Colorado as the secondary control state. These states were identified as balanced matches to California (the treated state) on baseline level and trend of cigarette sales and retail price using the itsamatch program in Stata [9]. We analyze the model using the new DDD-ITSA feature in the itsa program for Stata [7]. We specify regression with Newey–West standard errors [16] as the statistical model, and specify a single lag autoregressive structure (AR1) based on the results of the actest [17]. The Stata post-estimation command lincom is used to compute the Wald estimate for linear combinations of coefficients. The exact code used in this example is found in the Appendix.
@@ 3 Results
Table 5 presents the regression results for the DDD-ITSA model. First, we review how well the groups are balanced on baseline level and trend. $\beta_{4}$ and $\beta_{5}$ indicate that California and the primary control states (Idaho and Montana) are balanced on baseline level and trend ($P=0.347$ and $P=0.539$, respectively). A linear combination of terms of $(\beta_{4}-\beta_{8})$ and $(\beta_{5}-\beta_{9})$ indicate that California and Colorado (the secondary control state) are also balanced on baseline level and trend ($P=0.325$ and $P=0.711$ for level and trend, respectively; data not shown). Finally, $\beta_{8}$ and $\beta_{9}$ indicate that the two control groups are also balanced on baseline level and trend ($P=0.072$ and $P=0.880$, for level and trend, respectively).
Next, we evaluate treatment effects amongst the groups, with an emphasis on differences-in-differences in trends. $\beta_{7}$ indicates that California experienced a statistically significant annual net reduction (post-Proposition 99 – pre-Proposition 99) of 2.07-2.07 per-capita cigarette sales (in packs) as compared to the primary control states ($P=0.006$; 95% CI: 3.54-3.54, 0.61-0.61). Similarly, the linear combination of terms $(\beta_{7}-\beta_{11})$ indicates that California had a statistically significant annual net decrease of 1.76-1.76 per-capita cigarette sales (in packs) as compared to the secondary control ($P=0.020$; 95% CI: 3.24-3.24, 0.280-0.280). The $\beta_{11}$ coefficient indicates that the difference-in-differences between the primary and secondary control groups was a non-statistically significant annual decrease of 0.31-0.31 per-capita cigarette sales (in packs) ($P=0.723$; 95% CI: 2.03-2.03, $1.41$). As explained in the Methods section, the overall DDD-ITSA is the same as the treated unit versus the secondary control group, which in this example is a net annual reduction of 1.76-1.76 per-capita cigarette sales (in packs) ($P=0.020$; 95% CI: 3.24-3.24, 0.280-0.280). Figure 2 presents a graphic display of the results that are produced by itsa. Taken together, these results are similar to those presented in Figure 1(A), in which a treatment effect is realized against both sets of controls, and no false effects are noted between control groups. This example is intended to illustrate the implementation and interpretation of the DDD-ITSA framework under conditions where identifying assumptions appear plausible. In practice, however, violations of these assumptions (such as differential non-parallel trends or time-varying confounders that affect the component comparisons differently) may lead to biased estimates, underscoring the importance of careful assumption checking in applied analyses.
@@
Coefficient Label Estimate Std Err Z P 95% LCL 95% UCL
$\beta_{0}$ (const) $126.40$ $4.58$ $27.57$ $0.001$ $117.41$ $135.38$
$\beta_{1}$ (T) 1.43-1.43 $0.43$ 3.35-3.35 $0.001$ 2.27-2.27 0.59-0.59
$\beta_{2}$ (X) 11.43-11.43 $4.46$ 2.56-2.56 $0.010$ 20.17-20.17 2.68-2.68
[…]
$\beta_{4}$ (Z1) $5.83$ $6.20$ $0.94$ $0.347$ 6.33-6.33 $17.98$
[…]
$\beta_{8}$ (Z2) $12.29$ $6.84$ $1.80$ $0.072$ 1.11-1.11 $25.70$
[…]
  • Notes: Labels correspond to the coefficient labels in Equation (1); standard errors are Newey–West AR(1) adjusted.
@@ 4 Discussion
This paper formalizes the triple-difference interrupted time series (DDD-ITSA) design, an extension of the multiple-group ITSA framework that incorporates an additional control group to improve causal inference under a complementary set of identifying assumptions in health policy and intervention research. The DDD-ITSA approach combines the structural time-series logic of ITSA with the econometric rigor of difference-in-differences-in-differences estimators, offering researchers a powerful tool for estimating causal effects when standard two-group comparisons may remain confounded by unmeasured time-varying factors or differential trends across groups or comparison dimensions. The key contributions of this paper are the explicit articulation of the DDD-ITSA model, including clear parameter interpretation, and the worked example to illustrate the practical application of the DDD-ITSA design.
Researchers implementing DDD-ITSA face several issues with the design that require careful consideration. First, the choice of primary and secondary control groups is substantively important and should be guided by content expertise and empirical balance assessments. An additional consideration is the potential for spillovers or interference between groups. The validity of the DDD-ITSA design relies on the assumption that the intervention does not directly or indirectly affect outcomes in the control groups. In many policy settings, however, spillover effects may arise through mechanisms such as geographic proximity, behavioral responses, or market interactions. If such interference is present, the estimated treatment effects may be biased. Researchers should therefore carefully consider the plausibility of the no-interference assumption and, where possible, select control groups that are unlikely to be affected by the intervention. The primary control should be as similar as possible to the treatment unit on observed characteristics and pre-treatment trends [7, 9], while the secondary control provides an additional layer of adjustment, ideally one that shares the primary control’s susceptibility to common shocks but differs in ways that help isolate treatment effects.
@@
Third, researchers should carefully assess and account for autocorrelation, as failure to do so can lead to biased standard errors and invalid statistical inference, even when the underlying model is correctly specified. This can be accomplished by reviewing autocorrelation and partial autocorrelation functions [19] and formally testing for autocorrelation using appropriate diagnostics (e.g., the actest command in Stata [17]) and specifying error structures accordingly, whether through Newey–West standard errors [16], Prais–Winsten regression [20], Cochrane–Orcutt regression [21], instrumental variables [22], or any one of the autoregressive conditional heteroskedasticity family of estimators (see [23, 24] for a comprehensive discussion of econometric time series models).
Fourth, researchers should endeavor to collect as long of time series data as possible. Because the DDD-ITSA model includes multiple interaction terms, it requires substantially greater statistical resolution than SG-ITSA or MG-ITSA designs. In practice, this means that detecting meaningful effects, particularly changes in trends, typically requires a longer time horizon with sufficient pre- and post-intervention observations. Linden [18] found that in the MG-ITSA context, a longer study period provides the necessary statistical resolution to detect trend-based effects and mitigates the bias introduced by autocorrelation, which can otherwise severely undermine power. These considerations are likely to be even more important in DDD-ITSA, where additional parameters increase variance and may reduce precision in shorter series.
@@
The DDD-ITSA design is particularly well-suited to several common scenarios in health services and policy research. First, when researchers suspect that two-group comparisons may be confounded by concurrent policy changes, economic shocks, or secular trends that differentially affect groups or comparison dimensions, the addition of a second control group can difference out these common sources of bias. Second, DDD-ITSA is valuable when treatment effects are expected to vary across an additional dimension such as demographic groups or comparison dimensions, payer types, or clinical conditions. Third, the design is appropriate when researchers have access to multiple potential control groups but are uncertain which provides the most valid counterfactual. Rather than choosing arbitrarily or averaging across controls, the DDD-ITSA framework leverages both comparisons simultaneously, with the triple-difference estimand providing a summary effect that differences out bias common to the two component comparisons. This interpretation relies on the assumption that any bias affecting the underlying difference-in-differences comparisons operates similarly across them, rather than on the validity of any single comparison in isolation. Fourth, DDD-ITSA can serve as a robustness check for standard MG-ITSA findings. When two-group and three-group analyses yield consistent results, confidence in causal inferences increases. When results diverge, as in scenarios C and D in Figure 1, the pattern of findings can illuminate potential threats to validity and guide further investigation. For instance, a treatment effect in the secondary but not primary analysis (scenario C) might suggest spillover effects or confounding by events affecting the primary control, while effects in the primary but not secondary analysis (scenario D) might indicate that the secondary control is subject to unique influences that compromise its validity.
A key identifying assumption underlying the DDD-ITSA framework is that the difference between the two difference-in-differences comparisons would have remained constant in the absence of the intervention. While this is sometimes framed as requiring only a single parallel trends condition, in practice it implies that any bias present in the primary comparison must be replicated in the secondary comparison in order to be differenced out. If the factors generating non-parallel trends differ across these comparisons, the DDD estimate may remain biased. Researchers should therefore carefully consider the plausibility of this assumption when selecting control groups and interpreting and reporting results [25].
Despite the complexities of the design, the newly added DDD-ITSA functionality in the itsa package for Stata [7] simplifies the implementation and offers many options to ease interpretation and graphically display the outcomes (see the itsa help file for a complete description of options and additional examples). The interested reader will find all of the code used here for computing results of the Proposition 99 example in the Appendix.